OpenStudy (anonymous):

a rock at the end of the string follows a circular path if the string breaks, where does the rock go?

OpenStudy (anonymous):

Its tangent to the last position of the rock just before the string break.

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):



OpenStudy (anonymous):

i'm still really confused! help me please. use english :)

OpenStudy (anonymous):

In my previous picture, the circle is the path followed by the rock before string break.

OpenStudy (anonymous):

If the string break by the time i take the picture, the rock will follow stright line in the direction shown by arrow.

OpenStudy (anonymous):

 T -mg*cosdetha = 0 \[\frac{ m*v^2 }{ L } =mg*\cos \theta --->\cos \theta =\frac{ v^2 }{ L*g}\] v^2 -vo^2 =-2gR R =vo^2/2g